Kingbird Company’s record of transactions concerning part X for the month of April was as follows.

Purchases

Sales

April 1 (balance on hand) 210 @ $5.20 April 5 410
4 510 @ 5.30 12 310
11 410 @ 5.51 27 1,020
18 310 @ 5.56 28 150
26 710 @ 5.82
30 310 @ 6.03

Part 1

Correct answer iconYour answer is correct.

Calculate average-cost per unit. Assume that perpetual inventory records are kept in units only. (Round answer to 2 decimal places, e.g. 2.76.)

Average-cost per unit

$

eTextbook and Media

Attempts: 4 of 15 used

Part 2

New attempt is in progress. Some of the new entries may impact the last attempt grading.Your answer is incorrect.

Compute the inventory at April 30 on each of the following bases. Assume that perpetual inventory records are kept in units only. (1) First-in, first-out (FIFO). (2) Last-in, first-out (LIFO). (3) Average-cost. (Round final answers to 0 decimal places, e.g. $6,548.)
Answer #1

1. Calculation of average cost

Units rate purchase amount
210 $5.20 $1092
510 $5.30 $2703
410 $5.51 $2259.10
310 $5.56 $1723.60
710 $5.82 $4132.20
310 $6.03 $1869.30
Total $13779.20

Average cost per unit = $13779.20 / 2460 = $5.6 per unit

2. Total units = opening + purchase

= 2460 units

Units sold = 1890

Closing stock on April 30 = 2460 - 1890 = 570 units

FIFO Method - in this method units purchased first will be sold first , hence closing stock will consist recent purchases

= (310*$6.03) + (260*$5.82) = $3382.5

LIFO Method - In this method recently purchased goods will be sold first hence closing stock will contain opening stock and first purchase.

= (210*$5.20) + (360*$5.30)

= $3000

Average cost

= Closing stock * average cost

= 570 units * $5.6

= $3192